OpenStudy (anonymous):

Solve the equation 1/5(x-12)=1/4(x-15)

OpenStudy (anonymous):

its better to multiply whole equation by 100 5x + 8(100-x) = 560 5x + 800 - 8x = 560 3x = 800 - 560

OpenStudy (anonymous):

3x = 240 x = 80

OpenStudy (anonymous):

could you also assist with this question? 1/5(x-12)=1/4(x-15)

OpenStudy (anonymous):

\[1/5(x-12)=1/4(x-15) \]? this one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 1 }{ 5\left( x-12 \right) } - \frac{ 1 }{ 4\left( x-15 \right) } = 0\] \[\frac{ 4x-60-5x+60 }{ 20(x-12)(x-15) } = 0\] now you can solve ;)